OpenStudy (anonymous):

Harry has £2,200 He gives 1/4 to his son and 2/5 to his daughter How much does he keep for himself ? Can you show your working plz

OpenStudy (jamesj):

He gives away 1/4 + 2/5 = (5 + 8)/20 = 13/20 So he keeps 1 - 13/20 = (20-13)/20 = 7/20 Following so far?

OpenStudy (anonymous):

yep

OpenStudy (jamesj):

So in pounds, how much money does the father keep?

OpenStudy (anonymous):

he keeps 35% of 2200 which i forgot how to work out

OpenStudy (anonymous):

£770

OpenStudy (jamesj):

great

OpenStudy (anonymous):

THNX